An obstetrician wants to learn whether the amount of prenatal care and the wantedness of the pregnancy are associated. She randomly selects 939 women who had recently given birth and asks them to disclose whether their pregnancy was intended, unintended, or mistimed. In addition, they were to disclose when they started receiving prenatal care, if ever. The resulting contingency table is below. Are the requirements for performing a chi-square test of independence satisfied? Provide a rationale for you answer. wantedness_preg mos_preg_prior_care Crosstabulation mos preg prior_care > 5 mos. (or never) <3 mos.3 to 5 mos. 26 36.8 Total 652 41.7 652.0 83 5.3 83.0 204 13.0 204.0 939 60.0 939.0 wantedness preg intended Count 593 Expected Count 573.5 644.7 73.0 169 unintended Count її Expected Count mistimed Count 16 19 11.5 53 53.0 Expected Count 179. Total Count 826 826.0 60 Expected Count For the toolbar, press ALT+F10 (PC) or ALT+FN+F10 (Mac) TTT Arial ▼ 3(12pt) , T. -. -.剛 Click Save and Submit to save and submit. Click Save All Answers to save all answers Save All Answer O Type here to search

media%2F69d%2F69d1d344-94f1-4c72-a9db-8b

media%2F748%2F748c3d38-f016-4915-8ff3-8c
Answer #1

Observed (O) Prenatal Total >5 26 3 to 5 593 64 169 826 652 83 204 939 intended Wantedness unintende mistimed Total 19 53 16 60 Expected (E) Prenatal 3 to 5 >5 intended 573.538 36.8009 41.6613 Wantedness unintende 73.0117 4.68477 5.30351 mistimed 179.45 11.5144 13.0351 (O-E)A2/E Prenatal 3 to 5 >5 intended 0.66042 3.16999 1.80068 Wantedness unintende 1.1123 2.34606 6.11857 mistimed 0.60859 4.86649 0.67436

Chi-square=\sum \frac{ \left (O-E \right )^2}{E}=21.36

The P-value for Chi-square=21.36 for (3-1)(3-1)=4 degree of freedom id 0.0003.

Hence, we can conclude that there is an association relationship between the prenatal care and wantedness of the pregnancy at 0.05 level of significance.